# AI-driven View Guidance System in Intra-cardiac Echocardiography Imaging

### v1

@@ 5.4 State estimation result
As shown in Figure 2 (b), the proposed method can function as a current state estimator. For it to be an effective state estimator, it should provide a stable target state throughout the entire movement from the starting point to the target point. To validate this, we tested our method using a series of images representing movement from the starting point to the target point, visualizing the predicted fan direction for each case throughout the movement. We used four randomly selected subjects and chose a series of images ranging from a arbitrary starting state to a specific target volume.
@@
The results are shown in Figure 9. The colored triangles indicate the fan of the current image, while the black triangles represent the predicted fan. We determined the fan direction by measuring from the predicted position to the center point of the fan’s bottom long axis. This fan direction is separately visualized on the right side of each volume and fan image. Each standard deviation of the predicted position and the center point of the fan’s long axis is calculated and displayed next to the region, highlighted by a light gray elliptical shape.
@@
In Figure 9 (a), we assumed movement from the blue fan point to the RV target. From all intermediate images, we acquired the predicted fan direction, visualized as a black fan. The complete fan direction is displayed on the right side, showing that all predicted fan directions from intermediate images are consistently aligned. In Figure 9 (c), we assumed movement from the pink fan point to the RPV target. The predicted fan directions from all intermediate images are shown on the right side, demonstrating that the proposed method provides a stable target state for all intermediate images during movement and effectively functions as a state estimator.
@@
Figure 9: State estimation results. For several intermediate images during the movement from the start point to the target view, we acquired the target state from the proposed method and determined its fan direction. The left side displays the fan movement and its predicted position, while the right side shows the predicted fan direction vector. (a) RV target movement, (b) LAA target movement, (c) RPV target movement, and (d) LV target movement.
6 Conclusion and Discussion
@@
The intra-cardiac echocardiography (ICE) is an imaging tool used to capture images of the cardiac structures within the cardiac chamber. It is valuable for anatomical reviews, diagnosing congenital heart defects, or assessing valvular heart disease. Additionally, ICE is essential for interventional cardiology, guiding catheter ablation or the closure of septal defects. Despite its importance, effective manipulation of the ICE requires a skilled operator and substantial training due to the difficulty in achieving precise and repeatable controls. To address this issue, We proposed an AI-driven closed-loop view guidance system designed to simplify ICE manipulation based on the current ICE image. Specifically, there exist clinically predefined common views for most procedures, such as RV, LV, LPV, RPV, LAA, and ESO views. Our method provides the relative position and orientation vectors needed to guide the user to their preferred viewpoint from the current view, facilitating easier control of the ICE.
@@
To achieve this, we developed a Mamba-based regression model. The Mamba model, based on the State Space Model (SSM), outperforms traditional transformer architectures. Its selective scan and hardware-aware features enable efficient handling of long-range sequences. Our approach utilizes the MedMamba structure, which combines CNN and Mamba to efficiently capture both global and local features in medical images. For multi-target class regression, we incorporated a feature mixing layer to combine image features with target class codes, and a multi-head structure at the network’s end to separately estimate position and orientation. The network was trained using quantile loss to accurately estimate the boundaries of the predictions. We trained our regression model on a CARTO dataset containing view, position, orientation, and cardiac volume information.
The qualitative and quantitative results demonstrate that our method provides accurate target state information for view visualization. Whether the view is focused on specific structures or not, our method effectively supports view navigation. Additionally, when movement simulation test from the start point to the target state, the intermediate images stably show the target state well, proving that our method functions well as a state estimator.
@@
For future work, we plan to validate our system performance in a beating heart phantom. We believe that our method will simplify ICE manipulation for clinicians and enhance the overall workflow of the procedure.

### v3

@@ 5.5 Distance error for intermediate view while moving to the target view
When a user moves the ICE catheter from the current view to the target view based on the proposed method’s guidance, the system should continuously guide the user through all intermediate view images toward the target. To validate this, we tested our method using a sequence of images representing the movement from the starting point to the target point. The predicted fan direction was visualized at each step during the transition. For this evaluation, we selected four random subjects and utilized a series of images, each representing states ranging from an arbitrary starting point to a specific target volume.
@@
The results are shown in Figure 9. The colored triangles indicate the fan of the current image, while the black triangles represent the predicted fan. We determined the fan direction by measuring from the predicted position to the center point of the fan’s bottom long axis. This fan direction is separately visualized on the right side of each volume and fan image. Each standard deviation of the predicted position and the center point of the fan’s bottom long axis is calculated and displayed next to the region, highlighted by a light gray elliptical shape.
@@
In Figure 9 (a), we assumed movement from the blue fan point to the RV target. From all intermediate images, we acquired the predicted fan direction, visualized as a black fan. The complete fan direction is shown on the right side, showing that all predicted fan directions from intermediate images are consistently aligned. In Figure 9 (c), we assumed movement from the pink fan point to the RPV target. The predicted fan directions from all intermediate images are shown on the right side. It demonstrates that the proposed method provides a stable target state for all intermediate images during movement and effectively functions in a closed-loop system.
@@
Figure 9: Distance error for intermediate views while moving to the target view. For several intermediate images during the movement from the starting point to the target view, we acquired the target state from the proposed method and determined its fan direction. The left side shows the fan movement and its predicted position, while the right side shows the predicted fan direction vector. (a) RV target movement, (b) LAA target movement, (c) RPV target movement, and (d) LV target movement.
5.6 Human-in-the-loop validation
To validate our method in a realistic scenario, we conducted a discrete action guidance-based human-in-the-loop experiment. Since the method is designed and trained using real cardiac data, direct testing on human subjects was not feasible due to the inability to acquire real clinical imaging at arbitrary positions and orientations.
To address this limitation, we implemented a semi-simulation experiment using a real ICE catheter equipped with an EM sensor, as shown in Figure 10. The EM sensor served two purposes: 1) identifying the nearest real cardiac image from the CARTO dataset based on the ICE catheter’s current position and orientation, and 2) calculating errors relative to the target pose for performance validation. By leveraging the CARTO dataset, which contains pre-recorded clinical imaging, the system was able to map real clinical images to the ICE catheter’s pose at discrete intervals, simulating a realistic surgical environment. This approach, termed “semi-simulation”, combines real catheter manipulation with CARTO-derived imaging guidance. While the catheter is controlled in a physical setting, clinical imaging is mapped at discrete steps using CARTO data, compensating for the absence of direct clinical imaging acquisition. This hybrid approach bridges the gap between fully simulated setups and real-world clinical scenarios, providing a practical validation framework for the proposed system.
To further support the user’s adjustments, we incorporated angle markings on the ICE catheter to indicate the degree of knob rotation and bending. These markings allowed users to visually confirm the catheter’s movement during manipulation (shown in Figure 10). Although such angle-specific markings and mechanisms are not typically present in clinical environments, they were introduced as assumptions to establish a clear connection between the user’s manual adjustments and the proposed 4-DoF algorithm. This approach mimicked a real surgical scenario while leveraging CARTO data to provide realistic imaging guidance.
In more detail, the ICE catheter is initially placed in a straight condition as the home view ${\bf S}_{home}$. A target view (${\bf S}_{goal}$) is then randomly selected, and the proposed method provides guidance (${\bf J}_{home}^{goal}$) to assist the test engineers in manually manipulating the ICE catheter. The test engineers adjust the catheter based on the guidance provided on the display panel. When they believe the catheter has aligned with ${\bf J}_{home}^{goal}$, they confirm the position to the proposed system. If the catheter successfully reaches the target view (${\bf S}_{goal}$), the experiment concludes. Otherwise, if the moved state (${\bf S}_{curr}$) does not align with the target view, the proposed method iteratively generates new guidance (${\bf J}_{curr}^{goal}$) to direct further adjustments from the current view to the target view. This process is known as the re-planning step. All decisions and instructions are determined and provided by the proposed method. The procedure was performed on 10 cases from 10 different subjects, with three test engineers participating. The detailed steps of the procedure are shown below:
  1. 1.
    Place the ICE catheter in the straight condition as the home view ${\bf S}_{home}$ at a water chamber, and select the RV image of the randomly chosen subject from the real cardiac images in the CARTO testing dataset.
  2. 2.
    Choose randomly the target view (${\bf S}_{goal}$) from the 6 clinically defined views.
  3. 3.
    Proposed method displays (${\bf J}_{home}^{goal}$) on the display panel.
  4. 4.
    The test engineer manipulates the ICE catheter according to the guidance from the display panel.
  5. 5.
    The test engineer confirms to the proposed system once they have completed the manipulation of the ICE catheter to (${\bf S}_{goal}$)
  6. 6.
    If the proposed system confirms (${\bf S}_{curr}\approx{\bf S}_{goal}$), then proceed to step 7. Otherwise, during the re-planning phase, return to step 4 with updated guidance, (${\bf J}_{curr}^{goal}$), based on finding the nearest real cardiac images from the CARTO dataset using position and orientation data.
  7. 7.
    Calculate the error between the final state and the CARTO ground truth.
The representative results are shown in Figure 11. The blue fan represents the manually moved fan, the red fan indicates the target fan, and the gray fan shows the initial view. Figure 11 (a) shows a successful result after one movement, where the moved fan successfully crosses the target volume and aligns closely with the target fan. Across all cases, the average error between the final state and the real target value was $6.74[mm]$ and $(10.28,16.88,7.25)[deg]$ at step 7, confirming the applicability of our method in a real-world human-in-the-loop system.
Notably, we highlight the re-planning performance, where the error decreased after the re-planning step (Step 4-6). As shown in Figure 11 (b), the blue fan, representing the second moved fan, successfully crossed the target volume and aligned closely with the red target fan. The average error for the first movement was $15.12[mm]$ and $(12.02,17.49,14.41)[deg]$, which decreased to $4.30[mm]$ and $(6.98,10.94,3.78)[deg]$ after re-planning. These results demonstrate that the proposed method can effectively support continuous view navigation guidance through re-planning, ensuring precise movements and preventing missed alignments in a human-in-the-loop system.
Figure 10: Human-in-the-loop validation setup: The EM sensor was attached to the ICE catheter at the transducer location. The ICE catheter was controlled using the ICE controller, which includes knobs for adjusting the AP and LR directions, as well as for rotation and translation. For the experiment, the ICE catheter was positioned in a water chamber, as shown in the last column.
Figure 11: Human-in-the-loop validation results. The gray fan represents the initial view directed toward the RV. The green and blue fans denote the first and second manually adjusted fan views, respectively, while the red fan indicates the target view. (a) Direct movement from RV to LV. (b) Movement from RV to LPV through two re-planning steps.
6 Conclusion
@@
ICE is an imaging tool used to capture images of cardiac structures within the heart chambers. It is valuable for anatomical assessments, diagnosing congenital heart defects, and evaluating valvular heart disease. Additionally, ICE plays a crucial role in interventional cardiology, particularly in guiding catheter ablation and the closure of septal defects. Despite its importance, effective manipulation of the ICE requires a skilled operator and extensive training due to the difficulty in achieving precise and repeatable control. To address this challenge, we propose an AI-driven closed-loop view guidance system designed to simplify ICE manipulation based on the current ICE image. Specifically, there are clinically predefined common views for most procedures, such as RV, LV, LPV, RPV, LAA, and ESO views. Our method provides the relative position and orientation vectors necessary to guide the user to their preferred viewpoint from the current view, facilitating easier control of the ICE.
@@
To achieve this, we developed a Mamba-based regression model. The Mamba model, built upon the SSM, outperforms traditional transformer architectures. Its selective scan and hardware-aware features allow for efficient handling of long-range sequences. Our approach utilizes the MedMamba structure, which combines CNN and Mamba to effectively capture both global and local features in medical images. For multi-target class regression, we incorporated a feature mixing layer to combine image features with target class codes. Moreover, we added a multi-head structure at the network’s end to separately estimate position and orientation. The network was trained using quantile loss to accurately estimate the prediction boundaries. We trained our regression model on a CARTO dataset that includes view, position, orientation, and cardiac volume information.
The qualitative and quantitative results demonstrate that our method provides accurate target state information for view visualization. Whether the view focuses on specific structures or undefined regions, our method effectively supports view navigation. In particular, the human-in-the-loop validation emphasizes the feasibility of the proposed method in realistic scenarios. Although a gap between the real CARTO dataset and the EM sensor introduces some error, the overall results confirm the effectiveness and robustness of our method.
The human-in-the-loop validation demonstrates that our method can seamlessly integrate into real surgical procedures. To further enhance human-machine interaction, a speech or gesture recognition module can be incorporated, allowing for more intuitive control. Users can select from one of six representative views displayed on the interface panel, with view status updates provided at 50 [fps], matching the inference time of the method. The control panel displays real-time joint state information for four DoF and offers step-by-step guidance for adjusting each DoF. The system operates continuously until the user issues an end command.
@@
For future work, we plan to validate our system’s performance using a beating heart phantom. We believe that our method will simplify ICE manipulation for clinicians and enhance the overall workflow of the procedure.